# Fracton models from product codes

### 13 Dec 2023 v1

@@ Abstract
Figure 1: (a) The schematic depicts set inclusions for models representable by either product codes or the polynomial formalism [16, *Haah_Algebraic2016]. Explicit product code constructions of certain models are provided in the supplementary material [18]. (b,c) Classical Tanner graphs are shown, with bits represented by blue circles and checks by red squares. The HGP is represented graphically: in (b), a surface code is found as the HGP of two repetition codes, and in (c), the anisotropic $\mathbb{Z}_{2}$ Laplacian model [19] is obtained from a repetition code and a Laplacian code.
@@
Despite their different contexts, these topics are connected by certain individual results: for instance, Haah’s code, a Type-II fracton model, can be obtained via a generalization of the HGP known as the lifted product [21, 14]. One can also use classical codes to construct quantum “fractal spin liquids” [3], based on the polynomial formalism for translation invariant codes [16, *Haah_Algebraic2016] (Fig. 1).
@@
More broadly, an interesting observation suggests a fundamental connection between fracton order and HGP codes. To see this, we note that, for gapped phases of matter, the logarithm of the ground-state degeneracy determines the number of qubits that could in principle be encoded in the ground space. In conventional phases, this effective qubit number is independent of system size. This scaling is evidently incompatible with the number of logical qubits in typical HGP codes, which is polynomial in the system size. By contrast, such polynomial scaling is reminiscent of precisely the nontrivial dependence of the ground-state degeneracy on system size in fracton phases. Despite this tantalizing connection, a sharp understanding of the relationship between fracton order and product codes remains lacking.
@@
In this Letter, we provide such a framework, and demonstrate that product codes provide a natural setting for discovering new fracton phases of matter (Fig. 1). Our main results are threefold. First, we extend the definition of fracton order to stabilizer models without translation invariance or geometric locality. Based on this definition, we propose a set of criteria on the input classical codes of a product construction. When satisfied, the resulting quantum LDPC code realizes a nonlocal fracton phase. Second, we show that a recent model, purported to exhibit Type-I fracton order on an irregular graph, can be obtained as a hypergraph product code. Interestingly, while this particular case does not, in fact, constitute a fracton model, a generalized construction (utilizing a typical LDPC input code) does.
@@
Finally, we consider geometrically local product constructions, introducing a new family of classical LDPC codes defined on aperiodic tilings of the plane. Specifically, we consider a local two-dimensional code on the pinwheel tiling (Fig. 3), which we term the pinwheel code. We show that the HGP of the pinwheel code with a repetition code produces a local Type-I fracton model in three dimensions, and that the HGP of two pinwheel codes produces a local Type-II fracton model in four dimensions. These results open the door to the systematic discovery of new local fracton models using product codes. Somewhat intriguingly, the resulting models also exhibit evidence of a feature known as confinement (see discussions below), a favorable property of certain error correcting codes that is violated in most known fracton models [23, 24].
@@
Hypergraph product codes.—A classical LDPC code on $n$ bits with $m$ terms is represented by a $\kappa$-sparse $m\times n$ parity check matrix $H$ with elements in $\mathbb{F}_{2}=\{0,1\}$. It is characterized by a triplet $[n,k,d]$, where $k$ is the number of encoded logical bits and $d$ is the code distance [25]. The number of bits participating in each check is bounded above by a constant $\kappa_{c}$, and the number of checks acting on any given bit is bounded by $\kappa_{b}$; then $\kappa=\max(\kappa_{b},\kappa_{c})=O(1)$. Typical instances of LDPC codes are obtained by randomly sampling parity check matrices $H$ [18].
@@
Similarly, a CSS quantum code is represented by two parity check matrices $H_{X}$ and $H_{Z}$ specifying the action of the $X$ and $Z$ stabilizers, respectively [26, 27]. Both classical and quantum parity check matrices are maps taking a configuration of errors, represented by a $n$-dimensional vector over $\mathbb{F}_{2}$, to a syndrome or excitation, represented by an $m$-dimensional vector over $\mathbb{F}_{2}$.
@@
The hypergraph product (HGP) acts on a pair of classical “seed” codes, producing a CSS quantum code which is LDPC if the inputs are LDPC [22]. The quantum code resulting from the HGP of classical seed codes $H_{1}$ (with $n_{1}$ bits and $m_{1}$ checks) and $H_{2}$ (with $n_{2}$ bits and $m_{2}$ checks) is given by:
$\begin{gathered}H_{X}=\begin{bmatrix}H_{1}\otimes\mathbb{I}_{n_{2}}&\mathbb{I}_{m_{1}}\otimes H_{2}^{\top}\end{bmatrix},\\ H_{Z}=\begin{bmatrix}\mathbb{I}_{n_{1}}\otimes H_{2}&H_{1}^{\top}\otimes\mathbb{I}_{m_{2}}\end{bmatrix},\end{gathered}$ (1)
@@
$\begin{gathered}n_{q}=n_{1}n_{2}+m_{1}m_{2}~{},\\ k_{q}=k_{1}k_{2}+k_{1}^{\top}k_{2}^{\top}~{},\\ k_{X}^{\top}=k_{1}^{\top}k_{2}~{},~{}k_{Z}^{\top}=k_{1}k_{2}^{\top}~{},\\ d_{q}=\min(d_{1},d_{2},d_{1}^{\top},d_{2}^{\top})~{}.\end{gathered}$ (2)
Here, $k_{i}^{\top}$ and $d_{i}^{\top}$ denote the parameters of the transposed (or dual) code, with parity check matrix $H_{i}^{\top}$. The number of superselection sectors in the quantum code is given by $2^{k_{X}^{\top}+k_{Z}^{\top}}$ [18]. The prototypical example of a hypergraph product is the toric code, obtained from two cyclic repetition codes having $k_{1}=k_{1}^{\top}=k_{2}=k_{2}^{\top}=1$ and $d_{1}=n_{1}$, $d_{2}=n_{2}$ [22].
@@
Fracton order in nonlocal product codes.—Product constructions of stabilizer codes generically do not feature spatial symmetry or geometric locality; to connect to conventional fracton models [11, 12] and the local codes that we discuss later, we begin by providing a concrete definition of nonlocal fracton order. In particular, to ensure a well-defined thermodynamic limit, we sample from a random ensemble specifying a family of codes with connected Tanner graphs as $n_{q}\to\infty$  [^1] We have in mind codes which are far away from geometric locality, rather than hybrid geometries [55, 56].. Such a family is said to host fractons if: the logarithm of the ground state degeneracy scales polynomially in the system size  [^2] Note that this scaling is distinct from that of local fractons, which merely follows a polynomial envelope.; and excitations are pointlike  [^3] Finite-dimensional local fracton models with immobile loop excitations are known [57]. However such a case is not expected for the nonlocal models we consider. and have constrained mobility due to superselection sectors which also scale polynomially in the system size. An additional requirement in the case of nonlocal Type-II fractons is “confinement” [24], a natural consequence of nonlocality that implies the absence of string-like operators  [^4] The definition of confinement here is weaker than its usage elsewhere [23], as the energy cost is allowed to scale as any increasing function of the error weight, rather than strictly linearly.. The essence of confinement is that for sparse error patterns, additional errors always increase the number of syndromes.
Product constructions provide the opportunity to reason about quantum codes through investigation of their classical seeds. We propose three criteria for the classical seed codes that lead to nonlocal fracton order in the resulting HGP code: (i) rank deficiency, (ii) confinement, and (iii) isolability.
Rank deficiency refers to a property of the parity check matrix $H$. Since codewords are zero eigenvectors, a rank-deficient $H$ implies a large code space, with $k=n-\operatorname{\mathrm{rank}}(H)$. A rank-deficient seed code gives rise to superselection sectors in the resulting quantum code. A superselection sector is an equivalence class of syndromes under local errors; as the number of sectors grows, a diminishing number of excitation configurations may be dynamically connected by any physical process. This constitutes a strong mechanism for immobility, and is central to both local and nonlocal fracton models [33].
Intuitively, such an effect can be compared to the role of superselection sectors in, for example, the toric code, where they are labeled by elements of the anyon theory [34, 35]. Both cases identify dynamically disconnected excitation patterns; the difference is that topologically ordered phases like the toric code support only finitely many superselection sectors, whereas in an HGP code based on a rank-deficient seed code, the number of superselection sectors grows with the system size.
@@
Nonlocal Type-I fractons in HGP codes.—We consider hypergraph product codes, in which one classical seed is a cyclic repetition code and the other is either: case (i), a Laplacian code; or case (ii), a typical LDPC code (see [18] for details of each). This is motivated by a recently proposed stabilizer model for Type-I fracton order called the “anisotropic $\mathbb{Z}_{2}$ Laplacian model” [37]. While this model was originally derived from a series of results on dipole-conserving tensor gauge theories [38, 39, 19], it turns out to be precisely described by case (i) above [Fig. 1(c)] [18].
Figure 2: (a) Rank deficiency is exhibited for typical classical LDPC codes, from the linear scaling of $k$ with $n$; by contrast, the number of logical bits in a Laplacian code is on average bounded. (b) Confinement data is shown for both typical LDPC codes (blue) and Laplacian codes (red), on systems of $n=300$ and $500$ bits. Pictured is the minimum syndrome weight observed for randomly-sampled error configurations. Both constructions display evidence of confinement, as expected for nonlocal codes.
@@
Figure 3: (a) The pinwheel tiling substitution rule, taking one triangle to five, is shown. The graphs considered here are generated from two base tiles. (b,c) Graphs $G_{N}$ at generations $N=3,4$ are shown for the pinwheel tiling on a rectangular region. Each vertex hosts both a bit and a check, except for a fraction of boundary sites, whose checks are removed. The lowest-weight codeword is indicated in red for each. (d) The scaling of the number of logical bits $k\sim\sqrt{n}$ as well as the distance $d\sim n$ are shown for the pinwheel code, for several values $p=7,11,15$. Distances shown are exact, with the exception of the largest value having $k=65$, for which an upper bound is shown. (e) The minimum syndrome weight over sampled errors at $N=5$ provides numerical evidence for confinement in this local model. For low-weight errors, all error configurations are considered within balls of fixed radius centered on randomly sampled vertices. For high-weight errors, local truncations of logical operators illustrate the naive square-root upper bound. Though not exhaustive, such sampling can diagnose lack of confinement in practice [18].
@@
Local fracton models from aperiodic codes.— Though the stabilizer models discussed so far are only $\kappa$-local, it is natural to ask whether geometrically local HGP codes realize local fracton order. Whereas one well-known route to locality is to define a Tanner graph on a negatively curved manifold, which can embed expanders [41], we instead take a Euclidean approach by placing both bits and checks on the vertices of a two-dimensional aperiodic tiling.
@@
In contrast to typical LDPC codes, a $D$-dimensional local code cannot exhibit linear confinement. An upper bound is realized by the Ising model, wherein domain walls have polynomial energy cost with exponent $\frac{D-1}{D}$; more generally, in any local code, a codeword can be truncated within a $D$-ball at the cost of only a $(D-1)$-dimensional boundary syndrome. Such scaling is exhibited in the limit of large error weights in Fig. 3(e), and indeed follows a power law with exponent $1/2$. Though sampling is not guaranteed to capture the tails of the distribution, it is sufficiently thorough to distinguish the present cases from the Laplacian code on the square lattice, which, though rank deficient, is not confining [18]. We thus conjecture that the pinwheel code is confining, and saturates the upper bound.
Having demonstrated rank deficiency, confinement, and isolability, we now consider producing local fracton models via the hypergraph product of pinwheel codes. First, taking as inputs a pinwheel code and a cyclic repetition code yields a local Type-I fracton model in three dimensions. This code exhibits translation invariance only in the out-of-plane direction (i.e., perpendicular to the tiling). Second, taking the seed codes to be two instances of pinwheel codes yields a local Type-II fracton model in four dimensions, with parameters $k_{q}\sim n_{q}^{1/4}$, $d_{q}\sim\sqrt{n_{q}}$  [^7] While we have studied a particular rectangular patch under the substitution rules, more general regions are possible, and due to the lack of spatial symmetry non-identical codes will not have large coincident subgraphs.. We therefore largely capture the physics associated with the typical LDPC code [case (ii)] without having to appeal to nonlocal interactions or disorder. We conjecture that, similar to the four-dimensional Ising model, the Type-II pinwheel stabilizer model is confining without disorder, but, crucially, achieves this for point-like rather than extended excitations.

### 09 Apr 2024 v2

@@ Abstract
Figure 1: (a) Set inclusions are shown for stabilizer models. Some explicit product code constructions are provided in the supplementary material [16]. (b,c) Classical Tanner graphs are depicted, with bits represented by blue circles and checks by red squares. The HGP is represented graphically: in (b), a surface code is found as the HGP of two repetition codes, and in (c), the anisotropic $\mathbb{Z}_{2}$ Laplacian model [17, 18] is obtained from a repetition code and a Laplacian code.
@@
Despite their different contexts, these topics are connected by certain individual results: for instance, Haah’s code, a Type-II fracton model, can be obtained via a generalization of the HGP known as the lifted product [20, 14]. One can also use classical codes to construct quantum “fractal spin liquids” [3], based on the polynomial formalism for translation invariant codes [22, *Haah_Algebraic2016] (Fig. 1).
@@
More broadly, an interesting observation suggests a fundamental connection between fracton order and HGP codes. To see this, we note that, for gapped phases of matter, the logarithm of the ground-state degeneracy determines the number of qubits that could in principle be encoded in the ground space. In conventional phases, this effective qubit number is independent of system size. This scaling is evidently incompatible with the number of logical qubits in typical HGP codes, which is polynomial in the system size. By contrast, such polynomial scaling is reminiscent of precisely the nontrivial dependence of the ground-state degeneracy on system size in fracton phases. Nevertheless, a sharp understanding of the relationship between fracton order and product codes remains lacking.
@@
In this Letter, we provide such a framework, and demonstrate that product codes provide a natural setting for discovering new fracton phases of matter (Fig. 1). Our main results are threefold. First, we extend the definition of fracton order to stabilizer models without translation invariance or geometric locality. Based on this definition, we propose a set of criteria on the input classical codes of a product construction. When satisfied, the resulting quantum LDPC code realizes a nonlocal fracton phase. Second, we show that a recent model, purported to exhibit Type-I fracton order on an irregular graph, can be obtained as a HGP code. Interestingly, while this case does not, constitute a fracton model, a generalized construction (utilizing a typical LDPC input code) does.
@@
Finally, we consider geometrically local product constructions, introducing a new family of classical LDPC codes defined on aperiodic tilings of the plane. Specifically, we consider a local two-dimensional code on the pinwheel tiling (Fig. 3), which we term the pinwheel code. We show that the HGP of the pinwheel code with a repetition code produces a local Type-I fracton model in three dimensions, and that the HGP of two pinwheel codes produces a local Type-II fracton model in four dimensions. These results open the door to the systematic discovery of new local fracton models using product codes. The resulting models also exhibit evidence of a feature known as confinement (see discussions below), a favorable property the is lacking from most known fracton models [24, 25].
@@
Hypergraph product codes.—A classical LDPC code on $n$ bits with $m$ terms is represented by a $\kappa$-sparse $m\times n$ parity check matrix $H\in\mathbb{F}_{2}^{m\times n}$, where $\mathbb{F}_{2}=\{0,1\}$. It is characterized by a triplet $[n,k,d]$, where $k$ is the number of encoded logical bits and $d$ is the code distance [26]. The number of bits participating in each check is bounded above by a constant $\kappa_{c}$, and the number of checks acting on any given bit is bounded by $\kappa_{b}$; then $\kappa=\max(\kappa_{b},\kappa_{c})=O(1)$. Typical instances of LDPC codes are obtained by randomly sampling parity check matrices $H$ [16].
@@
Similarly, a CSS quantum code is represented by two parity check matrices $H_{X}$ and $H_{Z}$ specifying the action of the $X$ and $Z$ stabilizers, respectively [27, 28]. Both classical and quantum parity check matrices are maps taking a configuration of errors in $\mathbb{F}_{2}^{n}$, to a syndrome or excitation in $\mathbb{F}_{2}^{m}$.
@@
The HGP acts on a pair of classical “seed” codes, producing a CSS quantum code which is LDPC if the inputs are LDPC [21]. The quantum code resulting from the HGP of classical seed codes $H_{1}\in\mathbb{F}_{2}^{m_{1}\times n_{1}}$ and $H_{2}\in\mathbb{F}_{2}^{m_{2}\times n_{2}}$ is given by:
$H_{X}=\begin{bmatrix}H_{1}\otimes\mathbb{I}&\mathbb{I}\otimes H_{2}^{\top}\end{bmatrix},~{}H_{Z}=\begin{bmatrix}\mathbb{I}\otimes H_{2}&H_{1}^{\top}\otimes\mathbb{I}\end{bmatrix},$ (1)
@@
$\begin{matrix}n_{q}=n_{1}n_{2}+m_{1}m_{2}~{},~{}d_{q}=\min(d_{1},d_{2},d_{1}^{\top},d_{2}^{\top}),\\ k_{q}=k_{1}k_{2}+k_{1}^{\top}k_{2}^{\top}~{},~{}k_{X}^{\top}=k_{1}^{\top}k_{2}~{},~{}k_{Z}^{\top}=k_{1}k_{2}^{\top}~{}.\end{matrix}$ (2)
Here, $k_{i}^{\top}$ and $d_{i}^{\top}$ denote the parameters of the transposed (or dual) code, with parity check matrix $H_{i}^{\top}$. The number of superselection sectors in the quantum code is given by $2^{k_{X}^{\top}+k_{Z}^{\top}}$ [16]. The prototypical example of a HGP is the toric code, obtained from two cyclic repetition codes having $k_{1}=k_{1}^{\top}=k_{2}=k_{2}^{\top}=1$ and $d_{1}=n_{1}$, $d_{2}=n_{2}$ [21].
@@
Fracton order in nonlocal product codes.—Product constructions of stabilizer codes generically do not feature spatial symmetry or geometric locality; to connect to conventional fracton models [11, 12] and the local codes that we discuss later, we begin by providing a concrete definition of nonlocal fracton order. In particular, we sample from a random ensemble specifying a family of codes with connected Tanner graphs as $n_{q}\to\infty$. Such a family is said to host fractons if: the logarithm of the ground state degeneracy scales polynomially in the system size  [^1] Note that this scaling is distinct from that of local fractons, which merely follows a polynomial envelope.; and excitations are pointlike  [^2] While finite-dimensional local fracton models with immobile loop excitations are known [54], such a case is not expected for the nonlocal models we consider. and have constrained mobility due to superselection sectors which also scale polynomially in the system size. An additional requirement in the case of nonlocal Type-II fractons is “confinement” [25], a natural consequence of nonlocality that implies the absence of string-like operators  [^3] The definition of confinement here is weaker than its usage elsewhere [24], as the energy cost is allowed to scale as any increasing function of the error weight, rather than strictly linearly.. The essence of confinement is that for sparse error patterns, additional errors always increase syndrome weight.
Product constructions provide the opportunity to reason about quantum codes through investigation of classical seeds. We propose three criteria for classical codes leading to nonlocal fracton order in an HGP code: rank deficiency, confinement, and isolability.
Rank deficiency refers to a property of the parity check matrix $H$. Since codewords are zero eigenvectors, a rank-deficient $H$ implies a large code space, with $k=n-\operatorname{\mathrm{rank}}(H)$. A rank-deficient seed code gives rise to superselection sectors in the resulting quantum code. A superselection sector is an equivalence class of syndromes under local errors; as the number of sectors grows, a diminishing number of excitation configurations may be dynamically connected by any physical process  [^4] As in topological order, superselection sectors identify dynamically disconnected excitation patterns; the difference is that rather than supporting finitely many, labeled by elements of the anyon theory, in an HGP code based on a rank-deficient seed code, the number of superselection sectors depends on the system size.. This constitutes a strong mechanism for immobility, and is central to both local and nonlocal fracton models [34].
@@
Nonlocal Type-I fractons in HGP codes.—We consider HGP codes, in which one classical seed is a cyclic repetition code and the other is either: case (i), a Laplacian code; or case (ii), a typical LDPC code (see [16] for details of each). This is motivated by a recently proposed Type-I fracton model called the “anisotropic $\mathbb{Z}_{2}$ Laplacian model” [17, 18]. While originally derived from a series of results on dipole-conserving tensor gauge theories [36, 37, 38], this model is precisely described by case (i) above [Fig. 1(c)] [16].
Figure 2: (a) Rank deficiency is exhibited for typical classical LDPC codes, from the linear scaling of $k$ with $n$; by contrast, the number of logical bits in a Laplacian code is constant. (b) Confinement data is shown for both typical LDPC codes and Laplacian codes for $n=300$ and $500$. Pictured is the minimum syndrome weight of randomly sampled errors. Both constructions display evidence of confinement, as expected for nonlocal codes.
@@
Figure 3: (a) The pinwheel tiling substitution rule is shown. (b,c) Graphs $G_{N}$ at generations $N=3,4$ are shown for the pinwheel tiling on a rectangle. Each vertex hosts both a bit and a check, except for a fraction of boundary sites, whose checks are removed. The lowest-weight codeword is indicated in red. (d) The scaling of the number of logical bits $k\sim\sqrt{n}$ and distance $d\sim n$ are shown for the pinwheel code, for $p=7,11,15$. Distances shown are exact, except for the largest value $k=65$, for which an upper bound is shown. (e) The minimum syndrome weight over sampled errors at $N=5$ provides numerical evidence for confinement. For low-weight errors, all error configurations are considered within randomly sampled balls of fixed radius. For high-weight errors, truncated logical operators illustrate the naive square-root upper bound. Though not exhaustive, such sampling can diagnose lack of confinement in practice [16].
@@
Local fracton models from aperiodic codes.— Though the stabilizer models discussed so far are only $\kappa$-local, it is natural to ask whether geometrically local HGP codes realize local fracton order. Whereas one well-known route to locality is to define a Tanner graph on a negatively curved manifold, which can embed expanders [40], we instead take a Euclidean approach by placing both bits and checks on vertices of a two-dimensional aperiodic tiling.
@@
In contrast to typical LDPC codes, a $D$-dimensional local code cannot exhibit linear confinement. An upper bound is realized by the Ising model, wherein domain walls have polynomial energy cost with exponent $\frac{D-1}{D}$; more generally, in any local code, a codeword can be truncated within a $D$-ball at the cost of only a $(D-1)$-dimensional boundary syndrome. Such scaling $|s|\sim|e|^{1/2}$ is exhibited in the limit of large error weights in Fig. 3(e). Though sampling is not guaranteed to capture the tails of the distribution, it is sufficiently thorough to distinguish the present cases from the Laplacian code on the square lattice, which, though rank deficient, is not confining [16]. We thus conjecture that the pinwheel code is confining, and saturates the upper bound.
Having demonstrated rank deficiency, confinement, and isolability, we now consider the HGP of pinwheel codes as local fracton models. First, a pinwheel code and a cyclic repetition code yield a local Type-I fracton model in three dimensions, with translation symmetry perpendicular to the tiling only. Second, two instances of pinwheel codes yield a local Type-II fracton model in four dimensions, with parameters $k_{q}\sim n_{q}^{1/4}$, $d_{q}\sim\sqrt{n_{q}}$  [^7] While we have studied a particular rectangular patch under the substitution rules, more general regions are possible, and due to the lack of spatial symmetry non-identical codes will not have large coincident subgraphs.. We therefore largely capture the physics associated with the typical LDPC code [case (ii)] without appealing to nonlocal interactions or disorder. We conjecture that, similar to the four-dimensional toric code, the Type-II pinwheel stabilizer model is confining without disorder, but, crucially, achieves this for point-like rather than extended excitations.